OpenStudy (anonymous):

Martina's cow gives 1 3 over 4 milk each day. How much is this in quarts? Write your answer as a whole number or a mixed number in simplest form. Include the correct unit in your answer.

Directrix (directrix):

>>Martina's cow gives 1 3 over 4 milk each day. How much is this in quarts? If the question is meant to read; If Martina's cow gives 1 3/4 gallons of milk each day, how much is this in quarts?" 1 3/4 gallons = (4/4 + 3/4) gallons = 7/4 gallons. There are 4 quarts in one gallon. 7/4 times 4/1 = 7 quarts of milk.